OpenStudy (anonymous):

Help please? A teacher wrote expressions on cards and put them in a bag. Thomas and Kelly drew cards from the bag and added the expressions. Thomas drew the expressions 3x + 7 and 2x – 3. Kelly drew the expressions −6x + 3 and 14x + 2. How much greater was Kelly’s sum than Thomas’s sum?

OpenStudy (anonymous):

math

OpenStudy (anonymous):

what? @MyPhilHillName

TheSmartOne (thesmartone):

Ok, hello there! @ilovecrayons

OpenStudy (anonymous):

hi

TheSmartOne (thesmartone):

Let's first find the sum of Kelly's, shall we?

OpenStudy (anonymous):

okay

TheSmartOne (thesmartone):

-6x + 3 14x + 2 We need to add these two expressions -6x + 3 + 14x + 2 = ?? Can you add them? :)

OpenStudy (anonymous):

yes

TheSmartOne (thesmartone):

Ok, tell me what you get :)

OpenStudy (anonymous):

okay. thanks!

TheSmartOne (thesmartone):

remember, this isn't your final answer :)

OpenStudy (anonymous):

i got -8x+5

OpenStudy (anonymous):

is that right?

TheSmartOne (thesmartone):

kinda correct what is -6x + 14x = ? Hint: -6x + 14x = x(-6 + 14) = x(14-6) = ??

OpenStudy (anonymous):

okay, so x(14-6), no negatives? it is 8x?

TheSmartOne (thesmartone):

correct :)

OpenStudy (anonymous):

yay! okay, so how do I get the final answer?

TheSmartOne (thesmartone):

Now we need to find the sum of Thomas's expressions 3x + 7 2x - 3 3x + 7 + 2x - 3 = ??

OpenStudy (anonymous):

got it. okay. so 3x+2x-3= 5x + 4

OpenStudy (anonymous):

Is that right?

TheSmartOne (thesmartone):

correct :)

OpenStudy (anonymous):

great! what do i do now?

TheSmartOne (thesmartone):

Now we need to find the difference of Kelly's sum from Thomas's sum

TheSmartOne (thesmartone):

so \(\sf\Large 8x + 5 - (5x + 4) = 8x + 5 - 5x - 4 = ??\)

OpenStudy (anonymous):

is it 1+3x?

TheSmartOne (thesmartone):

and that's your answer :)

OpenStudy (anonymous):

yay! thanks so much!

TheSmartOne (thesmartone):

Great work, and thanks for cooperating! :D